OpenStudy (anonymous):

h^2-4h-21 polynomial expression, factoring of 21

OpenStudy (anonymous):

(h-7)(h+3) gives you -4 in the middle and -21 at the end.

OpenStudy (anonymous):

could you help me with two more

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

okay well here's the next one g^2-10g+21

OpenStudy (anonymous):

(g-7)(g-3) because it gives you -10 in the middle and the (-7)(-3) gives you the positive 21 at the end

OpenStudy (anonymous):

okay and my last one is n^2+10n+16

OpenStudy (anonymous):

(n+8)(n+2)

OpenStudy (anonymous):

do you know how to solve system of linear equations

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

omg ! great im horrible at them i need your help

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the problem is y=-3x-22 and underneath y=-8x-57

OpenStudy (anonymous):

all you do is equal the two equations together since both equal y so you get -3x-22=-8x-57 ==> add 57 to both sides -3x+35=-8x ==> add 3x to both sides 35=-5x ===> divide both sides by -5 x=-7 Plug -7 into both equations and you get y=-1 for both so x=-7 is the answer

OpenStudy (anonymous):

go it i got 2 more to do is it ok for you to help